OpenStudy (anonymous):

77= 9 over 5 * C +32. find C please help.

OpenStudy (anonymous):

I'm assuming it's 77 = 9 / (5*C + 32) cross multiply you get (5*C +32)*77 = 9 5*C + 2464 = 9 5*C = -2455 C = -491

OpenStudy (anonymous):

oops!!!

OpenStudy (dumbcow):

this is the conversion for Fahrenheit to celsius :) no its (9C/5) + 32

OpenStudy (anonymous):

ohhh good eye, i didn't see that and i made a really big mistake in my calculation 5C * 77 is not 5C

OpenStudy (anonymous):

so it's just 77 = (9/5)*C + 32 45 = (9/5)*C C = 25

OpenStudy (dumbcow):

no prob, i wouldve missed it too but i recognize the formula...people use parenthesis so we know whats what...thank you

OpenStudy (anonymous):

Thanks guys...